I need help with solving equations. I am wondering if I solved this problem correct. If anyone could let me know if I solved this problem correct and let me know, that would be great, thanks!
Problem:
0.6(x-50)=18-0.3(40-10x)
Answer that I came up with:
6x-300=18-120+30x
6x-300=-138+30x
6x-162=30x
-162=24x
-6.75=x

Answer by checkley71(8403)

.6(X-50)=18-.3(40-10X)
.6X-30=18-12+3X YOU FORGOT ABOUT THE DECIMAL POINT ON THE .6 & THE .3 ITEMS.
.6X-3X=18-12+30
-2.4X=36
X=36/-2.4
X=-15 ANSWER.
PROOF:
.6(-15-50)=18-.3(40-10*-15)
.6*-65=18-.3(40+150)
-39=18-.3*190
-39=18-57
-39=-39